# Tony Lindeberg

### v6

@@ 8 Summary and discussion
For this purpose, we have built the time-causal wavelet analysis on a special temporal smoothing kernel, denoted the time-causal limit kernel, and which constitutes the result of coupling an infinite set of truncated exponential kernels in cascade, with specially chosen time constants to obtain temporal scale covariance and self-similarity across scales. Specifically, we have in Sections 2 and 3 proposed to use temporal derivatives of this kernel as mother wavelets, with complementary normalization in a suitable $L_{p}$-norm. We have analyzed the continuous scaling properties of this representation, and also shown that the temporal derivatives of the time-causal limit kernel obey the admissibility condition for a mother wavelet. We have also argued that the temporal derivatives of the time-causal limit kernel can by theoretical arguments be regarded as canonical time-causal analogues of the derivatives of the regular Gaussian kernel, which for example, constitute the foundation for the Ricker wavelet.
@@
Then, we have in Section 4 turned to the problem of numerically approximating the time-causal limit kernel and its temporal derivatives for discrete signals. In this regard, first-order recursive filters constitute the natural analogue to the first-order integrators that equivalently describe the temporal smoothing with truncated exponential kernels, and which constitute the main computational primitives in the continuous theory. Thereby, we have describe a canonical way to compute discrete time-causal wavelet representations in Section 5, and also shown how this representation relates to bandpass representations, as can also be defined from continuous wavelets. Notably, the proposed implementation method is fully time-recursive, and does therefore lend itself directly to real-time processing on regular signal processing architectures. Specifically, as described in Algorithm 1, the temporal smoothing can be implemented in terms of a low number (4-8) first-order recursive filters coupled in cascade. The wavelet transforms are then obtained by applying compact temporal difference operators according to Equation (123) to the temporally smoothed data, complemented by complementary scale normalization. Thereby, real-time computations can be achieved on very low-capacity computational architectures.
@@
In summary, these experiments clearly demonstrate how the proposed family of time-causal wavelet representations can locally capture and reflect the temporal scales of locally dominant temporal structures based on a fully time-causal analysis of the input signals.
Based on these results, and the above mentioned principled theoretical foundations, we propose that this notion of time-causal wavelet analysis could be a valuable tool for signal processing tasks, where streams of signals are to be processed and analyzed in real time. Beyond real-time monitoring and time-series analysis and prediction, such applications may concern purposes of real-time feedback for e.g. control-loop systems, and as time-causal and scale-covariant temporal basis functions in learning-based approaches. Specifically, we argue that the proposed time-causal wavelet representations could be suitable for signals that may contain local variations over a rich span of temporal scales, and more generally for analysing physical or biophysical temporal phenomena, where a fully time-causal analysis is called for to be physically realistic.
Concerning further applications of the theory, we have intentionally in this presentation focused solely on theoretical and generic properties of the proposed families of time-causal wavelet representations, and not here aimed at developing more application-oriented experiments beyond the proof-of-concept studies in Sections 7 and 6.2. The motivation for this is that we want the results to be as timeless as possible, and not depending on specific properties of the experimental data, which may be different for different application domains, and may furthermore warrant the inclusion of additional more specialized computational mechanisms as well as integration with other signal processing approaches to build full-fledge applications.
@@
The raw zero-order (non-differentiated) time-causal limit kernel has additionally been successfully used as a replacement of the non-causal Gaussian window function in time-causal time-frequency analysis. For the corresponding notion of a time-causal and time-recursive analogue of Gabor filtering, corresponding to a time-causal analogue of the Morlet wavelet (Morlet et al. 1982a; 1982b), see Lindeberg (2025b). Compared to such time-causal time-frequency analysis, the time-causal wavelets developed in this paper require significantly less computations and lead to shorter temporal delays, with and with a better ability to handle rapid transients. For more general notions of a non-causal wavelets based on Hermite transforms, and which constitute extensions of the wavelets based on first- and second-order derivatives of the Gaussian kernel to higher orders of differentiation, see Markett (1993), Glaeske (2000) and Pandey and Phukan (2020).
In this treatment, we have, however, solely focused on temporal derivatives of orders 1 and 2, and leave the problem to future work of characterizing the properties of time-causal wavelets based on temporal derivatives of the time-causal limit kernel for higher orders of temporal differentiation.

### v9

@@ 8 Summary and discussion
For this purpose, we have built the time-causal wavelet analysis on a special temporal smoothing kernel, denoted the time-causal limit kernel, and which constitutes the result of coupling an infinite set of truncated exponential kernels in cascade, with specially chosen time constants to obtain temporal scale covariance and self-similarity across scales. Specifically, we have in Sections 2 and 3 proposed to use temporal derivatives of this kernel as mother wavelets, with complementary normalization in a suitable $L_{p}$-norm. We have analyzed the continuous scaling properties of this representation, and also shown that the temporal derivatives of the time-causal limit kernel obey the admissibility condition for a mother wavelet. We have also argued that the temporal derivatives of the time-causal limit kernel can by theoretical arguments be regarded as canonical time-causal analogues of the derivatives of the regular Gaussian kernel, which for example, constitute the foundation for the Ricker or Mexican hat wavelet.
@@
Then, we have in Section 4 turned to the problem of numerically approximating the time-causal limit kernel and its temporal derivatives for discrete signals. In this regard, first-order recursive filters constitute the natural analogue to the first-order integrators, that equivalently describe the temporal smoothing with truncated exponential kernels, and which constitute the main computational primitives in the continuous theory. Thereby, we have describe a canonical way to compute discrete time-causal wavelet representations in Section 5, and also shown how this representation relates to bandpass representations, as can also be defined from continuous wavelets. Notably, the proposed implementation method is fully time-recursive, and does therefore lend itself directly to real-time processing on regular signal processing architectures. Specifically, as described in Algorithm 1, the temporal smoothing can be implemented in terms of a low number (4-8) first-order recursive filters coupled in cascade. The wavelet transforms are then obtained by applying compact temporal difference operators according to Equation (123) to the temporally smoothed data, complemented by complementary scale normalization. Thereby, real-time computations can be achieved on very low-capacity computational architectures.
@@
In summary, these characterizations clearly demonstrate how the proposed family of time-causal wavelet representations can locally capture and reflect the temporal scales of locally dominant temporal structures based on a fully time-causal analysis of the input signals.
Based on these results, and the above mentioned principled theoretical foundations, we propose that this notion of time-causal wavelet analysis could be a valuable multi-purpose tool for signal processing tasks, where streams of signals are to be processed and analyzed in real time. Beyond real-time monitoring and time-series analysis and prediction, such applications may concern purposes of real-time feedback for e.g. control-loop systems, and as time-causal and scale-covariant temporal basis functions in learning-based approaches. Specifically, we argue that the proposed time-causal wavelet representations could be suitable for signals that may contain local variations over a rich span of temporal scales, and more generally for analysing physical or biophysical temporal phenomena, where a fully time-causal analysis is called for to be physically realistic, by not in any way attempting to access data from the future.
Concerning further applications of the theory, we have intentionally in this presentation focused mainly on theoretical and generic properties of the proposed families of time-causal wavelet representations, and not here aimed at developing more application-oriented experiments beyond the proof-of-concept studies in Sections 7 and 6.2. The motivation for this is that we want the results to be as timeless as possible, and not depending on specific properties of the experimental data, which may be different for different application domains, and may furthermore warrant the inclusion of additional more specialized computational mechanisms as well as integration with other signal processing approaches to build full-fledge applications.
@@
The raw zero-order (non-differentiated) time-causal limit kernel has additionally been successfully used as a replacement of the non-causal Gaussian window function in time-causal time-frequency analysis. For the corresponding notion of a time-causal and time-recursive analogue of Gabor filtering, corresponding to a time-causal analogue of the Morlet wavelet (Morlet et al. 1982a; 1982b), see Lindeberg (2025b). Compared to such time-causal time-frequency analysis, the time-causal wavelets developed in this paper require significantly less computations and lead to shorter temporal delays, and with a better ability to handle rapid transients. For more general notions of a non-causal wavelets based on Hermite transforms, and which constitute extensions of the wavelets based on first- and second-order derivatives of the Gaussian kernel to higher orders of differentiation, see Markett (1993), Glaeske (2000) and Pandey and Phukan (2020).
In this treatment, we have, however, solely focused on temporal derivatives of orders 1 and 2, and leave the topic to future work of characterizing the properties of time-causal wavelets based on temporal derivatives of the time-causal limit kernel for higher orders of temporal differentiation.